Kiesmel, Accounting, 6e Principles of Accoenting 2 Cus Season ticket s ter the Dr repced #3240 and ndde12hanegane. os An eouat neut er reven e s neva edat each gane n played when the seasonbeg nhe aroun 1826,000 ol the Uneanmed Ticket Revenue had been reoegnized as teverue How many season tickets did the Dingos se? Season tickets sold by Dingos 416,000 By the and of atober, How many home games had the Dingos slayed by the end of Oteber? Home games played by Dinges Prepare the of the Unearned Tiket Reversn fCendit accoune aes indented wh Crest Prepare tha
Answer #1

a) Season tickets sold by Dingos = 1416000/240 = 5900 tickets

b) Home games played by dingos =

Per game amount = 1416000/12 = 118000

home games played in october = 826000/118000 = 7 games

Journal entries

Date account and explanation debit credit
Cash 1416000
Unearned ticket revenue 1416000
(To record unearned revenue)
Unearned ticket revenue (1416000/12) 118000
Ticket revenue 118000
(to record adjusting entry)